OpenStudy (anonymous):

Solve each compound inequality: 0<-8b≤12

OpenStudy (anonymous):

Any ideas at all?

OpenStudy (anonymous):

What you want to do is, get the variable b on its own, so you have -8b, why not try dividing by -8 on both sides and see what you get :).

OpenStudy (gorv):

divide each term by -8

OpenStudy (gorv):

\[\frac{ 0 }{ -8}<\frac{ -8b }{ -8 }\le \frac{ 12 }{ -8 }\]

OpenStudy (gorv):

now try it